# Extrapolation for bilinear compact operators in the variable exponent setting

### v1

@@ 4. Factorization results for multilinear Muckenhoupt weights with variable exponents
One of the key ingredients for the proof of Theorem 1.1 is the following factorization result for the variable multilinear Muckenhoupt weights. The proof of this result is based on similar ideas of [18].
@@
Let $\gamma\in[0,m)$ be a fixed constant. Let $\vec{p}_{1}(\cdot),\vec{p}(\cdot)\in(\mathcal{E}^{0}_{(1,\infty)})^{m}$, $q_{1}(\cdot),q(\cdot)\in\mathrm{LH}\cap\mathscr{P}_{0}$ with
$\frac{1}{p_{1}(\cdot)}=\sum_{j=1}^{m}\frac{1}{p_{1,j}(\cdot)},\quad\frac{1}{p(\cdot)}=\sum_{j=1}^{m}\frac{1}{p_{j}(\cdot)},\quad\frac{1}{p_{1}(\cdot)}-\frac{1}{q_{1}(\cdot)}=\frac{1}{p(\cdot)}-\frac{1}{q(\cdot)}=\gamma.$
Let also $\vec{w}_{1}\in\mathcal{A}^{\gamma}_{\vec{p}_{1}(\cdot),q_{1}(\cdot)}$ and $\vec{w}\in\mathcal{A}^{\gamma}_{\vec{p}(\cdot),q(\cdot)}$. Given any $\theta\in(0,1)$, define $\vec{p}_{0}(\cdot)\in(\mathcal{P}_{0})^{m}$, $q_{0}(\cdot)\in\mathcal{P}_{0}$ by
@@
Then, there exist $\eta>0$, depending only on $n,m,$ $\vec{p}_{1}(\cdot),\vec{p}(\cdot),q_{1}(\cdot),q(\cdot)$ and the Muckenhoupt characteristics $\vec{w}_{1}]_{\mathcal{A}^{\gamma}_{\vec{p}_{1}(\cdot),q_{1}(\cdot)}}$, $\vec{w}]_{\mathcal{A}^{\gamma}_{\vec{p}(\cdot),q(\cdot)}},$ such that for all $\theta\in(0,\eta)$ all of the following hold:
@@
    $\vec{p}_{0}(\cdot)\in(\mathcal{E}^{0}_{(1,\infty)})^{m},$ $q_{0}(\cdot)\in\mathrm{LH}\cap\mathscr{P}_{0},$ and $\frac{1}{p_{0}(\cdot)}=\sum_{j=1}^{m}\frac{1}{p_{0,j}(\cdot)}$, $\frac{1}{p_{0}(\cdot)}-\frac{1}{q_{0}(\cdot)}=\gamma$.
@@
    $\vec{w}_{0}\in\mathcal{A}_{\vec{p}_{0}(\cdot),q_{0}(\cdot)}^{\gamma}.$
Proof.
Notice that once $\theta$ is fixed, $\vec{p}_{0}(\cdot),q_{0}(\cdot)$ and $\vec{w}_{0}(\cdot)$ are uniquely determined.
By the first part of the proof of [22, Lemma 3.1] we have that for all $j=1,\ldots,m$, by choosing
$\eta<\min\left(\frac{(p_{1,j})_{-}}{(p_{j})_{+}},1-\frac{1}{(p_{j})_{-}}\right),$
we can ensure that $p_{0,j}(\cdot)\in\mathcal{E}^{0}_{(1,\infty)}.$
Furthermore, it is easy to see that $mq(\cdot),mq_{1}(\cdot)\in\mathcal{E}^{0}_{(1,\infty)}$ and
$\frac{1}{mq(\cdot)}=\frac{1-\theta}{mq_{0}(\cdot)}+\frac{\theta}{mq_{1}(\cdot)},$
with $mq_{0}(\cdot)$ being completely determined from $\theta$, $mq_{1}(\cdot)$ and $mq(\cdot)$. Thus, by the same token, by choosing $\eta$ small enough depending only on $m$, $q_{1}(\cdot)$ and $q(\cdot)$ we can ensure that $mq_{0}(\cdot)\in\mathcal{E}^{0}_{(1,\infty)}$, so in particular $q_{0}(\cdot)\in\mathrm{LH}\cap\mathscr{P}_{0}.$
To sum up, by choosing $\eta$ small enough depending only on $\vec{p}_{1}(\cdot),\vec{p}(\cdot),q_{1}(\cdot),q(\cdot)$ we can ensure that $\vec{p}_{0}(\cdot)\in(\mathcal{E}^{0}_{(1,\infty)})^{m}$ and $q_{0}(\cdot)\in\mathrm{LH}\cap\mathscr{P}_{0}.$ If by chance $(q_{1})_{-},q_{-}>1,$ then by picking $\eta$ even smaller we can ensure in addition $(q_{0})_{-}>1.$
It is trivial to verify the relation $\frac{1}{p_{0}(\cdot)}-\frac{1}{q_{0}(\cdot)}=\gamma.$
To prove that for small enough $\eta$, $\vec{w}_{0}\in\mathcal{A}_{\vec{p}_{0}(\cdot),q_{0}(\cdot)}^{\gamma}$, it suffices by Lemma 3.4 to show that
$\begin{cases}w_{0,j}^{-\frac{1}{m}}\in\mathcal{A}_{mp_{0,j}^{\prime}(\cdot)},\quad j=1,\ldots,m\\ \nu_{\vec{w}_{0}}^{\frac{1}{m}}\in\mathcal{A}_{mq_{0}(\cdot)}.\end{cases}$
Since $\vec{w}_{1}\in\mathcal{A}_{\vec{p}_{1}(\cdot),q_{1}(\cdot)}^{\gamma}$ and $\vec{w}\in\mathcal{A}_{\vec{p}(\cdot),q(\cdot)}^{\gamma}$, by Lemma 3.4 we have
$\begin{cases}w_{1,j}^{-\frac{1}{m}}\in\mathcal{A}_{mp_{1,j}^{\prime}(\cdot)},\quad j=1,\ldots,m\\ \nu_{\vec{w}_{1}}^{\frac{1}{m}}\in\mathcal{A}_{mq_{1}(\cdot)}.\end{cases}$
@@
$\begin{cases}w_{j}^{-\frac{1}{m}}\in\mathcal{A}_{mp_{j}^{\prime}(\cdot)},\quad j=1,\ldots,m\\ \nu_{\vec{w}}^{\frac{1}{m}}\in\mathcal{A}_{mq(\cdot)}.\end{cases}$
Moreover, by the proof of [6, Lemma 1.2] we have that the associated variable exponent Muckenhoupt characteristics are all bounded from above by $\vec{w}_{1}]_{\mathcal{A}^{\gamma}_{\vec{p}_{1}(\cdot),q_{1}(\cdot)}}^{\frac{1}{m}},$ respectively $\vec{w}]_{\mathcal{A}^{\gamma}_{\vec{p}(\cdot),q(\cdot)}}^{\frac{1}{m}},$ up to implied constants depending only on $n,m,$ as well as $\vec{p}_{1}(\cdot)$ and $q_{1}(\cdot)$, respectively $\vec{p}(\cdot)$ and $q(\cdot).$
Fix now $j=1,\ldots,m.$ Set $v:=w_{j}^{-\frac{1}{m}}$ and $r(\cdot):=mp_{j}^{\prime}(\cdot),$ extending these notations analogously to their endpoint counterparts. Observe that $r(\cdot),r_{1}(\cdot)\in\mathcal{E}^{0}_{(1,\infty)}$, $v_{1}\in\mathcal{A}_{r_{1}(\cdot)},$ $v\in\mathcal{A}_{r(\cdot)},$
$v=w_{j}^{-\frac{1}{m}}=(w_{0,j}^{-\frac{1}{m}})^{1-\theta}(w_{1,j}^{-\frac{1}{m}})^{\theta}=v_{0}^{1-\theta}v_{1}^{\theta},$
@@
$\frac{1}{r(\cdot)}$ $=\frac{1}{mp_{j}(\cdot)^{\prime}}=\frac{1}{m}\left(1-\frac{1}{p_{j}(\cdot)}\right)=\frac{1}{m}\left(1-\theta+\theta-\frac{1-\theta}{p_{0,j}(\cdot)}-\frac{\theta}{p_{1,j}(\cdot)}\right)$
$=\frac{1-\theta}{mp_{0,j}^{\prime}(\cdot)}+\frac{\theta}{mp_{1,j}^{\prime}(\cdot)}=\frac{1-\theta}{r_{0}(\cdot)}+\frac{\theta}{r_{1}(\cdot)}.$
Thus, a direct application of [22, Lemma 3.1] in the diagonal setting and the named variable exponents and weights shows that we can pick $\eta$ small enough, depending only on $n,$ $r_{1}(\cdot),r(\cdot),$ $v_{1}]_{\mathcal{A}_{r_{1}(\cdot)}}$ and $v]_{\mathcal{A}_{r(\cdot)}}$, so that for any $\theta<\eta$ one has $v_{0}\in\mathcal{A}_{r_{0}(\cdot)}.$ Note that here we only need a somewhat simplified version of the second part of the proof of [22, Lemma 3.1]: by considering only the diagonal case, one avoids having to use [11, Proposition 5.4], so that a direct application of the reverse Hölder inequality, Lemma 3.3 is possible.
Similarly, one can ensure that $\nu_{\vec{w}_{0}}^{\frac{1}{m}}\in\mathcal{A}_{mq_{0}(\cdot)}.$ This concludes the proof. ∎
In view of later applications of our extrapolation results, we also need a slightly varied version of Lemma 4.1. Even though the changes are restricted essentially to the notational level, we include a precise statement and a short explanation for the sake of clarity.
@@
Let $t$ be a fixed constant with
$1<t<\min\left(\min\{(p_{1,j})_{-}:~j=1,\ldots,m\},\min\{(p_{j})_{-}:~j=1,\ldots,m\}\right),$
and let $\gamma\in\left[0,\frac{m}{t}\right)$ be a fixed constant. Consider $\vec{p}_{1}(\cdot),\vec{p}(\cdot)\in(\mathcal{E}^{0}_{(1,\infty)})^{m}$ and $q_{1}(\cdot),q(\cdot)\in\mathrm{LH}\cap\mathscr{P}_{0}$ with
$\frac{1}{p_{1}(\cdot)}=\sum_{j=1}^{m}\frac{1}{p_{1,j}(\cdot)},\quad\frac{1}{p(\cdot)}=\sum_{j=1}^{m}\frac{1}{p_{j}(\cdot)},\quad\frac{1}{p_{1}(\cdot)}-\frac{1}{q_{1}(\cdot)}=\frac{1}{p(\cdot)}-\frac{1}{q(\cdot)}=\gamma.$
Let also $\vec{w}_{1}\in\mathcal{A}^{\gamma}_{\vec{p}_{1}(\cdot),q_{1}(\cdot)}$ and $\vec{w}\in\mathcal{A}^{\gamma}_{\vec{p}(\cdot),q(\cdot)}$ such that $\vec{w}_{1}^{t}\in\mathcal{A}^{t\gamma}_{\frac{\vec{p}_{1}(\cdot)}{t},\frac{q_{1}(\cdot)}{t}}$ and $\vec{w}^{t}\in\mathcal{A}^{t\gamma}_{\frac{\vec{p}(\cdot)}{t},\frac{q(\cdot)}{t}}$.
Given any $\theta\in(0,1)$, define $\vec{p}_{0}(\cdot)\in(\mathcal{P}_{0})^{m}$, $q_{0}(\cdot)\in\mathcal{P}_{0}$ by
@@
Then, there exists $\eta>0$, depending only on $n,m,$ $\vec{p}_{1}(\cdot),\vec{p}(\cdot),q_{1}(\cdot),q(\cdot),$ $t$ and the Muckenhoupt characteristics $\vec{w}_{1}]_{\mathcal{A}^{\gamma}_{\vec{p}_{1}(\cdot),q_{1}(\cdot)}}$, $\vec{w}]_{\mathcal{A}^{\gamma}_{\vec{p}(\cdot),q(\cdot)}},$ such that for all $\theta\in(0,\eta)$ all of the following hold:
@@
    $\vec{p}_{0}(\cdot)\in(\mathcal{E}^{0}_{(1,\infty)})^{m},$ $q_{0}(\cdot)\in\mathrm{LH}\cap\mathscr{P}_{0},$ and $\frac{1}{p_{0}(\cdot)}=\sum_{j=1}^{m}\frac{1}{p_{0,j}(\cdot)}$, $\frac{1}{p_{0}(\cdot)}-\frac{1}{q_{0}(\cdot)}=\gamma.$
@@
    $\vec{w}_{0}\in\mathcal{A}_{\vec{p}_{0}(\cdot),q_{0}(\cdot)}^{\gamma}.$
  5. (5)
    $\vec{w}_{0}^{t}\in\mathcal{A}_{\frac{\vec{p}_{0}(\cdot)}{t},\frac{q_{0}(\cdot)}{t}}^{t\gamma}$.
@@
Set $\delta:=t\gamma\in[0,m)$ and define the $\pi\text{-}\kappa\text{-}\delta\text{-}\omega$” counterparts
@@
and extend these notation analogously to their counterparts for the endpoint 1. Then, we clearly have
$\vec{\pi}_{1}(\cdot),\vec{\pi}(\cdot)\in(\mathcal{E}^{0}_{(1,\infty)})^{m},$
$\kappa_{1}(\cdot),\kappa(\cdot)\in\mathrm{LH}\cap\mathscr{P}_{0}$
$\vec{\omega}_{1}\in\mathcal{A}^{\delta}_{\vec{\pi}_{1}(\cdot),\kappa_{1}(\cdot)},$
$\vec{\omega}\in\mathcal{A}^{\delta}_{\vec{\pi}(\cdot),\kappa(\cdot)}.$
Moreover, it holds
@@
The $\pi\text{-}\kappa\text{-}\delta\text{-}\omega$” counterparts for the endpoint 0 are defined directly in terms of the already defined $\pi\text{-}\kappa\text{-}\delta\text{-}\omega$” coounterparts by
@@
Thus, by applying Lemma 4.1 both for the original objects as well as their $\pi\text{-}\kappa\text{-}\delta\text{-}\omega$” counterparts, we can achieve that by picking $\eta$ small enough, depending on the same quantities as in Lemma 4.1 and in addition $t$, the relation $\theta<\eta$ already ensures that
  • $\vec{p}_{0}(\cdot)\in(\mathcal{E}^{0}_{(1,\infty)})^{m},$ $q_{0}(\cdot)\in\mathrm{LH}\cap\mathscr{P}_{0},$ $\frac{1}{p_{0}(\cdot)}-\frac{1}{q_{0}(\cdot)}=\gamma,$ $\vec{w}_{0}\in\mathcal{A}_{\vec{p}_{0}(\cdot),q_{0}(\cdot)}^{\gamma}$, and in addition $(q_{1})_{-},q_{-}>1\Rightarrow(q_{0})_{-}>1,$
  • $\vec{\pi}_{0}(\cdot)\in(\mathcal{E}^{0}_{(1,\infty)})^{m},$ $\kappa_{0}(\cdot)\in\mathrm{LH}\cap\mathscr{P}_{0},$ $\frac{1}{\pi_{0}(\cdot)}-\frac{1}{\kappa_{0}(\cdot)}=\delta,$ $\vec{\omega}_{0}\in\mathcal{A}_{\vec{\pi}_{0}(\cdot),\kappa_{0}(\cdot)}^{\delta}$.
Therefore, it now remains to check that our definition of the $\pi\text{-}\kappa\text{-}\delta\text{-}\omega$” counterparts for the endpoint 0 automatically yields the desired $t$-scaling, that is that we indeed have

### v2

@@ 4. Factorization results for multilinear Muckenhoupt weights with variable exponents
One of the key ingredients for the proof of Theorem 1.1 is the following factorization result for the variable multilinear Muckenhoupt weights. The proof of this result is based on similar ideas of [undefq, undefr, undefv].
@@
Let $(\vec{p}(\cdot),q(\cdot),\vec{r},s)$ and $(\vec{p}_{1}(\cdot),q_{1}(\cdot),\vec{r},s)$ be proper $m$-admissible quadruples such that
$\frac{1}{p_{1}(\cdot)}-\frac{1}{q_{1}(\cdot)}=\frac{1}{p(\cdot)}-\frac{1}{q(\cdot)}=\gamma\in[0,\infty).$
Moreover, let
$\vec{w}\in\mathcal{A}_{(\vec{p}(\cdot),q(\cdot)),(\vec{r},s)}$
and
$\vec{w}_{1}\in\mathcal{A}_{(\vec{p}_{1}(\cdot),q_{1}(\cdot)),(\vec{r},s)}.$
Given any $\theta\in(0,1)$, define $\vec{p}_{0}(\cdot)\in(\mathscr{P}_{0})^{m}$, $q_{0}(\cdot)\in\mathscr{P}_{0}$ by
@@
Then, there exists $\eta>0$ such that for all $\theta\in(0,\eta)$ all of the following hold:
@@
    $(\vec{p}_{0}(\cdot),q_{0}(\cdot),\vec{r},s)$ is a proper $m$-admissible quadruple with
    $\quad\frac{1}{p_{0}(\cdot)}-\frac{1}{q_{0}(\cdot)}=\gamma.$
  2. (2)
    There holds
    $\vec{w}_{0}\in\mathcal{A}_{(\vec{p}_{0}(\cdot),q_{0}(\cdot)),(\vec{r},s)}.$
@@
The proof relies on [undefv, Lemma 3.20]. Since the notation used there is slightly different from the one used here (see (2) of Remark 2.8 for a comparison), we restate [undefv, Lemma 3.20] for the reader’s convenience in our notation.
4.2 Lemma ([undefv, Lemma 3.20]).
Let $(p(\cdot),q(\cdot),r,s)$ and $(p_{1}(\cdot),q_{1}(\cdot),r,s)$ be proper $1$-admissible quadruples such that $r\geq 1$ and
$\frac{1}{p_{1}(\cdot)}-\frac{1}{q_{1}(\cdot)}=\frac{1}{p(\cdot)}-\frac{1}{q(\cdot)}=\gamma\in[0,1).$
Let also
$w\in\mathcal{A}_{(p(\cdot),q(\cdot)),(r,s)}$
and
$\vec{w}_{1}\in\mathcal{A}_{(p_{1}(\cdot),q_{1}(\cdot)),(r,s)}.$
Given any $\theta\in(0,1)$, define $p_{0}(\cdot)\in\mathscr{P}_{0}$, $q_{0}(\cdot)\in\mathscr{P}_{0}$ by
$\frac{1}{p(\cdot)}$ $=\frac{1-\theta}{p_{0}(\cdot)}+\frac{\theta}{p_{1}(\cdot)},$
as well as the weight $w_{0}$ by
$w$ $=w_{0}^{1-\theta}w_{1}^{\theta}.$
Then, there exists $\eta>0$, such that for all $\theta\in(0,\eta)$ all of the following hold:
@@
    $(p_{0}(\cdot),q_{0}(\cdot),r,s)$ is a proper $1$-admissible quadruple with $\frac{1}{p_{0}(\cdot)}-\frac{1}{q_{0}(\cdot)}=\gamma$.
  2. (2)
    There holds
    $w_{0}\in\mathcal{A}_{(p_{0}(\cdot),q_{0}(\cdot)),(r,s)}.$
It might seem at a first glance that in the way we have rephrased [undefv, Lemma 3.20] and we have missed assumptions on the ranges of the exponents stipulated in [undefv, Lemma 3.20]. However, as explained in (2) of Remark 2.8, this is not the case.
Now we prove Lemma 4.1.
Proof of Lemma 4.1.
Notice that once $\theta$ is fixed, $\vec{p}_{0}(\cdot),q_{0}(\cdot)$ and $\vec{w}_{0}(\cdot)$ are uniquely determined. It is also trivial to verify that
$\frac{1}{p_{0}(\cdot)}-\frac{1}{q_{0}(\cdot)}=\gamma.$
Step 1. Fix $j\in\{1,\ldots,m\}$. Pick $a\in(0,\infty)$ with $ar_{j}\geq 1$. Then, the quadruples
$(ap_{j}(\cdot),ap_{j}(\cdot),ar_{j},m\sigma_{j})\quad\text{and}\quad(ap_{1,j}(\cdot),ap_{1,j}(\cdot),ar_{j},a\sigma_{j})$
are proper $1$-admissible with $ar_{j}\geq 1$ and
$\frac{1}{ap_{j}(\cdot)}-\frac{1}{ap_{j}(\cdot)}=\frac{1}{ap_{1,j}(\cdot)}-\frac{1}{ap_{1,j}(\cdot)}=0\in[0,1).$
Moreover, in view of our assumptions, by combining Lemma 3.5 with Lemma 3.8, we deduce that
$w_{j}^{\frac{1}{a}}\in\mathcal{A}_{ap_{j}(\cdot),(ar_{j},a\sigma_{j})}$
@@
$w_{1,j}^{\frac{1}{a}}\in\mathcal{A}_{ap_{1,j}(\cdot),(ar_{j},a\sigma_{j})}.$
Clearly
$w_{j}^{\frac{1}{a}}=(w_{0,j}^{\frac{1}{a}})^{1-\theta}(w_{1,j}^{\frac{1}{a}})^{\theta}.$
Thus, by applying Lemma 4.2, we deduce that there exists $\eta_{j}>0$, so that for all $\theta\in(0,\eta_{j})$, $(ap_{0,j}(\cdot),ap_{0,j}(\cdot),ar_{j},a\sigma_{j})$ is a proper 1-admissible quadruple and $w_{0}^{\frac{1}{a}}\in\mathcal{A}_{ap_{0,j}(\cdot),(ar_{j},a\sigma_{j})}$. By Lemma 3.8, this means $w_{0}\in\mathcal{A}_{p_{0,j}(\cdot),(r_{j},\sigma_{j})}$. Moreover, we have $ap_{0,j}(\cdot)\in\mathrm{LH}$, therefore also $p_{0,j}(\cdot)\in\mathrm{LH}$.
Step 2. Pick $b\in(0,\infty)$ with
$b>\max\left\{\gamma,\frac{1}{r}\right\}.$
Then, the quadruples
$(bp(\cdot),bq(\cdot),br,bs)\quad\text{and}\quad(bp_{1}(\cdot),bq_{1}(\cdot),br,bs)$
are proper $1$-admissible with $br\geq 1$ and
$\frac{1}{bp(\cdot)}-\frac{1}{bq(\cdot)}=\frac{1}{bp_{1}(\cdot)}-\frac{1}{bq_{1}(\cdot)}=\frac{\gamma}{b}\in[0,1).$
Moreover, in view of our assumptions, by combining Lemma 3.5 with Lemma 3.8, we deduce that
$\nu_{\vec{w}}^{\frac{1}{b}}\in\mathcal{A}_{(bp(\cdot),bq(\cdot)),(br,bs)}$
@@
$\nu_{\vec{w}_{1}}^{\frac{1}{b}}\in\mathcal{A}_{(bp_{1}(\cdot),bq_{1}(\cdot)),(br,bs)}.$
Clearly
$\nu_{\vec{w}}^{\frac{1}{b}}=(\nu_{\vec{w}_{0}}^{\frac{1}{b}})^{1-\theta}(\nu_{\vec{w}_{1}}^{\frac{1}{b}})^{\theta}.$
Thus, by applying Lemma 4.2, we deduce that there exists $\eta_{\nu}>0$, so that for all $\theta\in(0,\eta_{\nu})$, $(bp_{0}(\cdot),bq_{0}(\cdot),br,bs)$ is a proper 1-admissible quadruple and $\nu_{\vec{w}_{0}}^{\frac{1}{b}}\in\mathcal{A}_{(bp_{0}(\cdot),bq_{0}(\cdot)),(br,bs)}$. By Lemma 3.8, this means $\nu_{\vec{w}_{0}}\in\mathcal{A}_{(p_{0}(\cdot),q_{0}(\cdot)),(r,s)}$. Moreover, we have $bp_{0}(\cdot),bq_{0}(\cdot)\in\mathrm{LH}$, therefore also $p_{0}(\cdot),q_{0}(\cdot)\in\mathrm{LH}$.
Step 3. Take $\eta:=\min\{\eta_{1},\ldots,\eta_{m},\eta_{\nu}\}$. Then, the quadruple $(\vec{p}_{0}(\cdot),q_{0}(\cdot),\vec{r},s)$ is proper $m$-admissible. Moreover, we have
$w_{0}\in\mathcal{A}_{p_{0,j}(\cdot),(r_{j},\sigma_{j})},\quad\forall j=1,\ldots,m$
and
$\nu_{\vec{w}_{0}}\in\mathcal{A}_{(p_{0}(\cdot),q_{0}(\cdot)),(r,s)}.$
By Lemma 3.5 we deduce that $\vec{w}_{0}\in\mathcal{A}_{(\vec{p}_{0}(\cdot),q_{0}(\cdot)),(\vec{r},s)}$.
Step 4. If $(q_{1})_{-},q_{-}>1$, then by the proof of [undefv, Lemma 3.1] we have that by picking $\eta$ even smaller we can ensure in addition that $(q_{0})_{-}>1$. ∎
In view of later applications of our extrapolation results, we also need a varied version of a special case of Lemma 4.1. Even though the changes are restricted essentially to the notational level, we include a precise statement and a short explanation for the sake of clarity.
@@
Let $(\vec{p}(\cdot),q(\cdot),\vec{1},\infty)$ and $(\vec{p}_{1}(\cdot),q_{1}(\cdot),\vec{1},\infty)$ be proper $m$-admissible quadruples with
$\frac{1}{p(\cdot)}-\frac{1}{q(\cdot)}=\frac{1}{p_{1}(\cdot)}-\frac{1}{q_{1}(\cdot)}=\gamma\in[0,\infty).$
Let $t\in(0,\infty)$ be a fixed constant with
$t<\min\{(p_{j})_{-}:~j=1,\ldots,m\}$
and
$t<\min\{(p_{1,j})_{-}:~j=1,\ldots,m\}.$
Let
$\vec{w}\in\mathcal{A}_{\vec{p}(\cdot),q(\cdot)}$
and
$\vec{w}_{1}\in\mathcal{A}_{\vec{p}_{1}(\cdot),q_{1}(\cdot)}$
such that in addition one has
$\vec{w}^{t}\in\mathcal{A}_{\frac{\vec{p}(\cdot)}{t},\frac{q(\cdot)}{t}}$
and
$\vec{w}_{1}\in\mathcal{A}_{\frac{\vec{p}_{1}(\cdot)}{t},\frac{q_{1}(\cdot)}{t}}.$
Given any $\theta\in(0,1)$, define $\vec{p}_{0}(\cdot)\in(\mathscr{P}_{0})^{m}$, $q_{0}(\cdot)\in\mathscr{P}_{0}$ by
@@
Then, there exists $\eta>0$ such that for all $\theta\in(0,\eta)$ all of the following hold:
@@
    $(\vec{p}_{0}(\cdot),q_{0}(\cdot),\vec{1},\infty)$ is a proper $m$-admissible quadruple with $\frac{1}{p_{0}(\cdot)}-\frac{1}{q_{0}(\cdot)}=\gamma$.
  2. (2)
    $\vec{w}_{0}\in\mathcal{A}_{\vec{p}_{0}(\cdot),q_{0}(\cdot)}$.
@@
    $\vec{w}_{0}^{t}\in\mathcal{A}_{\frac{\vec{p}_{0}(\cdot)}{t},\frac{q_{0}(\cdot)}{t}}$.
@@
By a direct application of Lemma 4.1 we have that items (1), (2) and (3) are all satisfied if we choose $\eta$ small enough.
Set now $\delta:=t\gamma\in[0,\infty)$ and define the $\pi\text{-}\kappa\text{-}\delta\text{-}\omega$” counterparts
@@
and extend these pieces of notation analogously to their counterparts for the endpoint 1. Then, we clearly have that
$(\vec{\pi}(\cdot),\kappa(\cdot),\vec{1},\infty)\quad\text{and}\quad(\vec{\pi}_{1}(\cdot),\kappa_{1}(\cdot),\vec{1},\infty)$
are proper $m$-admissible quadruples with
@@
Moreover, by assumption it holds that
$\vec{\omega}_{1}\in\mathcal{A}_{\vec{\pi}_{1}(\cdot),\kappa_{1}(\cdot)},$
$\vec{\omega}\in\mathcal{A}_{\vec{\pi}(\cdot),\kappa(\cdot)}.$
The $\pi\text{-}\kappa\text{-}\delta\text{-}\omega$” counterparts for the endpoint 0 are defined directly in terms of the already defined $\pi\text{-}\kappa\text{-}\delta\text{-}\omega$” counterparts by
@@
Thus, by applying Lemma 4.1 for the $\pi\text{-}\kappa\text{-}\delta\text{-}\omega$” counterparts, we can achieve that by picking $\eta$ small enough, the relation $\theta<\eta$ ensures that
  • $(\vec{\pi}_{0}(\cdot),\kappa_{0}(\cdot),\vec{1},\infty)$ is a proper $m$-admissible quadruple with $\frac{1}{\pi_{0}(\cdot)}-\frac{1}{\kappa_{0}(\cdot)}=\delta$, and
  • $\vec{\omega}_{0}\in\mathcal{A}_{\vec{\pi}_{0}(\cdot),\kappa_{0}(\cdot)}$.
Therefore, it only remains to check that our definition of the $\pi\text{-}\kappa\text{-}\delta\text{-}\omega$” counterparts for the endpoint 0 automatically yields the desired $t$-scaling, that is that we indeed have